# Non-Abelian Statistics for Bosonic Symmetry-Protected Topological Phases

### v3

@@ III.2 Braiding schemes with tri-junctions
depending on the presence or absence of the local defect. Here, $\left|{\xi(T)}\right\rangle\equiv U(T)\left|{\xi}\right\rangle$ and $\left|{\eta(T)}\right\rangle\equiv U(T)\left|{\eta}\right\rangle$. In conclusion, the braiding matrix is $\sigma^{z}$ for conventional braiding and $i\sigma^{y}$ for defect-assisted braiding in the new bases $\{\left|{\xi}\right\rangle,\left|{\eta}\right\rangle\}$. To further demonstrate the topological nature of the two classes of braiding statistics, in Appendix C, we rigorously compute the non-Abelian Berry phases in the tri-junction setting. In particular, we show that the two classes of statistics fall into distinct topological classes, the classification of which can be related to the Stiefel-Whitney classes of real vector bundles. The defect at the junction acts as a singular point, and leads to a topological Berry phase $e^{i\Delta\phi}=\langle\beta_{2}|U(T)|\beta_{1}\rangle\langle\beta_{1}|U(T)|\beta_{2}\rangle$, with $\Delta\phi=\pi$ for the braiding. Appendix D presents a more intuitive derivation of the braiding results for more general and larger tri-junctions in the dimerized limit.
Figure 3: Numerical time evolution of zero-mode wavefunctions for two braiding schemes. (a) Amplitudes of and (b) phase difference between $\left|{\beta_{1}(t)}\right\rangle$ and $\left|{\beta_{2}(t)}\right\rangle$ for the first type of braiding scheme. (c) Amplitudes and (d) phase difference for the second type of braiding scheme. The horizontal axis represents the evolution time in units of the total braiding time $T$. The strong and weak couplings take $v_{\text{max}}=1$ and $v_{\text{min}}=0.1$. The simulations use a precision of time step $\Delta t=0.1$ and an elementary step time $T_{0}=60$ for the first scheme ($T_{0}=36$ for the second). Each braiding consists of 15 elementary steps of zero-mode motion, giving a total braiding time $T=15T_{0}=900$ for the first scheme ($T=15T_{0}=540$ for the second). The red horizontal line denotes the theoretical prediction for $\Delta\phi=\mathrm{Arg}(\langle\beta_{2}|\beta_{1}(t)\rangle)-\mathrm{Arg}(\langle\beta_{1}|\beta_{2}(t)\rangle)\,\mathrm{mod}\,2\pi$ after braiding. The vertical dashed line marks the end of the braiding, after which the system is kept static to illustrate the stabilized value of $\Delta\phi$, as shown in the figure.
@@
$|\overline{00}\rangle=|\emptyset\rangle|\xi\rangle,\,|\overline{10}\rangle=|\xi\rangle|\emptyset\rangle,\,|\overline{01}\rangle=|\emptyset\rangle|\eta\rangle,\,|\overline{11}\rangle=|\eta\rangle|\emptyset\rangle.$
@@ IV Numerical simulations
We present in this section the numerical results to support the theory in the previous sections. We demonstrate the time evolution of ground-state wavefunctions and compute the average fidelities of the braiding matrices for both conventional and defect-assisted braidings. In particular, we focus on the half-filled sector $\mathcal{H}_{0}$ spanned by $\left|{\beta_{1}}\right\rangle$ and $\left|{\beta_{2}}\right\rangle$, where the non-Abelian nature is prominently manifested.
We consider specific tri-junctions with $16$ sites depicted in Figs. 1(c) and 1(d), which implement the first and the second types of braiding scheme, respectively. For the first type of braiding scheme, the dynamics of the system are governed by the time-dependent Hamiltonian
@@
The Hamiltonians in Eqs. (39) and (40) are referred to as the first and the second type of tri-junction Hamiltonians, respectively. Initially, the coupling strengths are set according to the configurations in Figs. 1(c) and 1(d), with the strong and weak couplings $v=1$ and $v^{\prime}=0.1$. To investigate the effect of system size on gap protection for tri-junctions, we numerically calculate the energy gap $\delta E$ between the ground state and the first excited state of the first type of tri-junction Hamiltonian at $t=0$ as a function of tri-junction size, shown in Fig. 1(f). The gap $\delta E$ is slightly larger for smaller sizes and rapidly approaches a constant value as the size increases. A similar behavior is observed for the second type of tri-junction Hamiltonian, and is thus omitted. For $t>0$, the energy gap $\delta E$ deviates slightly from a constant value and oscillates periodically with time. Nevertheless, the amplitude of the oscillation is very small, allowing the gap to be approximated as constant during the braiding process. Therefore, the proposed tri-junction configuration provides gap protection throughout the braiding process.
@@
To implement the braiding, we employ the scheme in Fig. 1(e) to manipulate zero modes. In each step, we take three sites, decrease the initial strong coupling strength $v$ and increase the weak coupling strength $v^{\prime}$ over a time period $T_{0}$, following the functional dependence in Eq. (24), with $v_{\text{max}}=v$ and $v_{\text{min}}=v^{\prime}$. This adiabatic tuning shifts the zero mode by two sites. The braiding process consists of $15$ moving steps for both types of scheme. For the second type of braiding scheme, special care is needed when the zero mode passes near the local defect. The subtlety is to ensure that the defect remains isolated both before and after the zero mode passes through it. To achieve this, we first decouple the zero mode from the chain as it moves adjacent to the defect, ensuring that it is localized on a single site. We then exchange the zero mode with the local defect in a dimerized-limit fashion. Finally, we recouple the zero mode to the chain and complete the remaining movements. This approach guarantees that the local defect returns to its original state after braiding. The critical aspect of our method is maintaining defect isolation by decoupling the zero mode when it passes through. In Appendix E, we quantitatively analyze errors that arise when defect isolation is not maintained, confirming that our scheme effectively prevents such errors.
@@
Figure 4: Numerical time evolution of zero-mode wavefunctions under random hopping disorders in the dimerized tri-junctions. (a) and (b) show the evolution of the amplitudes of $\left|{\beta_{1}(t)}\right\rangle$, $\left|{\beta_{2}(t)}\right\rangle$, and their phase difference under complex-valued random hopping disorders. (c) and (d) show the evolution under real-valued random hopping disorders. The calculations are performed on the dimerized tri-junctions with $10$ sites [Fig. 7(b)].The horizontal axis shows the evolution time in units of the total braiding time $T$. The strong and weak couplings take $v_{\text{max}}=1$ and $v_{\text{min}}=0$ Simulations are performed with a time step $\Delta t=0.1$ and an elementary step time $T_{0}=80$. Each braiding consists of $9$ elementary steps of zero-mode motion, giving a total time $T=9T_{0}=720$. The red horizontal line represents the theoretical prediction of $\Delta\phi=\text{Arg}(\langle\beta_{2}|\beta_{1}(t)\rangle)-\text{Arg}(\langle\beta_{1}|\beta_{2}(t)\rangle)\,\text{mod}\,2\pi$ in the absence of dynamical symmetry breaking. The vertical dashed line marks the end of the braiding process, after which the system is held static to demonstrate the stabilized value of $\Delta\phi$. The values of $\Delta\phi$ for both disorder cases are indicated in the figure, and the numerical results are averaged over $100$ random disorder realizations. The disorder strength is introduced as a fluctuation of the exchange coupling, $\delta v_{i}=s_{i}v_{i}$, where $s_{i}$ follows a Gaussian distribution with zero mean and standard deviation $\sigma=0.05$.
@@
$U(T)\approx\text{lim}_{\Delta t\rightarrow 0}e^{-iH(T)\Delta t}e^{-iH(T-\Delta t)\Delta t}\cdots e^{-iH(0)\Delta t}.$
@@
To further quantify the accuracy of the numerical results, we calculate the average fidelity $F$ of the braiding matrix (see Appendix F for the definition). For defect-assisted braiding, the average fidelity is $F(U,i\sigma^{y})=0.99999999$, while for conventional braiding, it is $F(U,\sigma^{z})=0.999$. The decrease in fidelity for conventional braiding is due to the small energy splitting $\Delta E$ between two ground states caused by finite-size effects [Fig. 1(b)]. In contrast, defect-assisted braiding is less affected by the finite-size effect. These results show that there is no need to tune the system to the dimerized limit to achieve high fidelity. The scheme exhibits substantial robustness against finite-size effects.
Figure 5: Illustration of the experimental scheme. (a) The minimal model with four spins. Spins 22 and 33 collectively serve as a defect. Spins 11 and 44 support zero modes in the dimerized limit. The dashed lines represent spin-exchange couplings. (b) and (c) Initial configurations for the first and second type of braiding schemes. A red oval encircling two spins represents the bell state $\frac{1}{\sqrt{2}}(\left|{10}\right\rangle-\left|{01}\right\rangle)$. (d) The time-dependent coupling strengths as functions of time to achieve braiding process. The coupling strength is in unit of its maximum value $v$. The ascending (descending) slope takes the form $\chi(t/T_{0})$ ($1-\chi(t/T_{0})$). (e) Tunable qubit-qubit interaction via a coupler, where $\omega_{q}$ and $\omega_{c}$ are the qubit and coupler frequencies, respectively, and $g_{1}$, $g_{2}$ and $g_{12}$ are fixed nearest neighbour and next nearest neighbour couplings. (f) Superconducting qubit tri-junction with tunable qubit-qubit interactions. Red and blue lines indicate the direct qubit-qubit coupling and qubit-coupler coupling, respectively.
Finally we comment the analogy nature of the simulation of the quasiparticles with finite-size topological systems. The energy gap between the ground state and first excited state can be numerically obtained as a function of tri-junction size [see Fig. 1(f)]. As determined by the coupling strengths, the gap decreases slightly versus system size and approaches a constant quickly as the size increases. This tells that the gap in a finite system is nearly identical to that in the thermodynamic limit. In consequence, the SPNA statistics for the topological phases defined in the thermodynamic limit can be achieved in the finite-size system as protected by the same gap. In other words, the increase of the system size shall also not significantly increase the complexity of the simulation of the SPNA statistics, marking a key difference from the digital simulation. Finally, thanks to the gap protection, the experimental realization of the SPNA statistics can be achieved with the minimal tri-junction configuration, a strategy that we adopt in Sec. VI.
@@ V Symmetry protection mechanism of the braiding statistics
Now we turn to investigate in detail the symmetry protection mechanism of SPNA statistics for bosonic SPT phases. From numerical results we show that the dynamical symmetry breaking may be induced by complex random disorders, but not by real ones.
For convenience, we consider the second type of braiding scheme in the dimerized limit (dynamical symmetry breaking for the first type of braiding scheme is similar). We introduce random imaginary hopping terms:
@@
to the Hamiltonian of the junction (Eq. (74)), where $\delta v_{i}(t)=s_{i}v_{i}(t)$ and each $s_{i}$ is a Gaussian random variable with zero mean and standard deviation $\sigma=0.05$. This term preserves anti-unitary symmetry generated by $C^{A}$ and the particle number conservation, but breaks the unitary particle-hole symmetry. In Figs. 4(a) and 4(b), we plot the evolution of amplitudes of zero-mode states $\left|{\beta_{1}}\right\rangle$, $\left|{\beta_{2}}\right\rangle$ and their phase difference under the Hamiltonian $H(t)+H_{p}(t)$. Fig. 4(a) shows that $|\langle\beta_{1}|\beta_{2}(t=T)\rangle|$ and $|\langle\beta_{2}|\beta_{1}(t=T)\rangle|$ reach $0.996$, while $|\langle\beta_{1}|\beta_{1}(t=T)\rangle|$ and $|\langle\beta_{2}|\beta_{2}(t=T)\rangle|$ are 0, indicating that the states run out of the subspace $\mathcal{H}_{0}$. Fig. 4(b) shows that the phase difference $\Delta\phi$ stabilizes at the value deviating from $\pi$ after braiding, confirming that the non-unitary symmetry undergoes dynamical breaking, causing the braiding statistics to become ill-defined in the ground-state subspace. In comparison, Figs. 4(c) and 4(d) show the evolution under the Hamiltonian $H(t)-iH_{p}(t)$, which represents real random hopping disorders and preserves all the unitary and anti-unitary symmetries. These results show that braiding statistics remains robust under real random hopping disorders. This highlights the crucial role of unitary symmetries in maintaining the robustness of SPNA statistics in bosonic SPT phases and the resilience of the system against real hopping disorders.
@@ VI Experimental schemes
To determine the experimental parameters, we calculate the average fidelity as a function of the elementary moving step time $T_{0}$ [Fig. 6]. The fidelity calculations indicate that the dominate factor limiting the performance of the scheme is the adiabaticity of braiding process, further supporting the conclusion that the system is gap-protected. Moreover, the saturation of the fidelity at large $T_{0}$ further confirms the topological nature of both classes of braiding statistics, since dynamical phases would vary with time and thus could not produce a stable plateau. The two classes of statistics originate from non-Abelian Berry phases in distinct topological classes, as shown in Appendix C and confirmed by numerical tests demonstrating robustness against local modifications of the evolution functions in Fig. 5(d). From Fig. 6, We find that $\frac{T_{0}v}{\hbar}>25$ achieves average fidelities $F>0.99$, where $v$ represents the maximum spin-exchange coupling strength during the braiding. This implies that the total braiding time $T=3T_{0}>\frac{75\hbar}{v}$. To reduce the time required for the braiding operation, it is essential to make the spin-exchange coupling sufficiently large. With typical superconducting qubit coupling strengths reaching up to $20$ MHz Salathé et al. (2015), the minimum braiding time is $3.75\mu s$ for a fidelity over $0.99$. The decoherence time $T_{2}$ for superconducting qubits is around $2\sim 10\mu s$ Bylander et al. (2011). We also note that implementing experiments with superconducting qubits by applying the non-adiabatic shortcut methods may further reduce the braiding time.

### v4

@@ III.2 Braiding schemes with tri-junctions
in the case with or without the local defect. Here, $\left|{\xi(T)}\right\rangle\equiv U(T)\left|{\xi}\right\rangle$ and $\left|{\eta(T)}\right\rangle\equiv U(T)\left|{\eta}\right\rangle$. In this new single-qubit basis $\{\left|{\xi}\right\rangle,\left|{\eta}\right\rangle\}$, we achieve the non–commutative braiding matrices $\sigma^{z}$ for conventional braiding and $i\sigma^{y}$ for defect-assisted braiding.
Figure 3: Numerical time evolution of zero-mode wavefunctions for two braiding schemes. (a) Amplitudes of and (b) phase difference between $\left|{\beta_{1}(t)}\right\rangle$ and $\left|{\beta_{2}(t)}\right\rangle$ for the first type of braiding scheme (conventional braiding). (c) Amplitudes and (d) phase difference for the second type of braiding scheme (defect-assisted braiding). The horizontal axis represents the evolution time in units of the total braiding time $T$. The strong and weak couplings take $v_{\text{max}}=1$ and $v_{\text{min}}=0.1$. The simulations use a precision of time step $\Delta t=0.1$ and an elementary step time $T_{0}=60$ for the first scheme ($T_{0}=36$ for the second). Each braiding consists of 15 elementary steps of zero-mode motion, giving a total braiding time $T=15T_{0}=900$ for the first scheme ($T=15T_{0}=540$ for the second). The red horizontal line denotes the theoretical prediction for $\Delta\phi=\mathrm{Arg}(\langle\beta_{2}|\beta_{1}(t)\rangle)-\mathrm{Arg}(\langle\beta_{1}|\beta_{2}(t)\rangle)\,\mathrm{mod}\,2\pi$ after braiding. The vertical dashed line marks the end of the braiding, after which the system is kept static to illustrate the stabilized value of $\Delta\phi$, as shown in the figure.
@@
$|\overline{00}\rangle=|\emptyset\rangle|\xi\rangle,\,|\overline{10}\rangle=|\xi\rangle|\emptyset\rangle,\,|\overline{01}\rangle=|\emptyset\rangle|\eta\rangle,\,|\overline{11}\rangle=|\eta\rangle|\emptyset\rangle.$ (38)
@@ IV Numerical simulations
We present in this section the numerical results to support the rigorous theory in the previous section. We demonstrate the time evolution of ground-state wavefunctions and compute the average fidelities of the braiding matrices for both conventional and defect-assisted braidings. In particular, we focus on the half-filled sector $\mathcal{H}_{0}$ spanned by $\left|{\beta_{1}}\right\rangle$ and $\left|{\beta_{2}}\right\rangle$, where the non-Abelian nature is prominently manifested.
We consider specific tri-junctions with $16$ sites depicted in Figs. 1(c) and 1(d), which implement the first and second types of braiding scheme, respectively. For the first type of braiding scheme, the dynamics of the system are governed by the time-dependent Hamiltonian
@@
The Hamiltonians in Eqs. (41) and (42) are referred to as the first and the second type of tri-junction Hamiltonians, respectively. Initially, the coupling strengths are set according to the configurations in Figs. 1(c) and 1(d), with the strong and weak couplings $v=1$ and $v^{\prime}=0.1$. To investigate the effect of system size on gap protection for tri-junctions, we numerically calculate the energy gap $\delta E$ between the ground state and the first excited state of the first type of tri-junction Hamiltonian at $t=0$ as a function of tri-junction size, shown in Fig. 1(f). The gap $\delta E$ is slightly larger for smaller sizes and rapidly approaches a constant value as the size increases. A similar behavior is observed for the second type of tri-junction Hamiltonian. For $t>0$, the energy gap $\delta E$ deviates slightly from a constant value and oscillates periodically with time. Nevertheless, the amplitude of the oscillation is very small, allowing the gap to be approximated as constant during the braiding process. Therefore, the proposed tri-junction configuration provides gap protection throughout the braiding process.
@@
To implement the braiding, we employ the scheme in Fig. 1(e) to manipulate zero modes. In each step, we take three sites, decrease the initial strong coupling strength $v$ and increase the weak coupling strength $v^{\prime}$ over a time period $T_{0}$, following the functional dependence in Eq. (25), with $v_{\text{max}}=v$ and $v_{\text{min}}=v^{\prime}$. This adiabatic tuning shifts the zero mode by two sites. The braiding process consists of $15$ moving steps. For the second type of braiding scheme, special care is needed when the zero mode passes near the local defect. The subtlety is to ensure that the defect remains isolated both before and after the zero mode passes through it. To achieve this, we first decouple the zero mode from the chain as it moves adjacent to the defect, ensuring that it is localized on a single site. We then exchange the zero mode with the local defect in a dimerized-limit fashion. Finally, we recouple the zero mode to the chain and complete the remaining movements. This approach guarantees that the local defect returns to its original state after braiding. The critical aspect of our method is maintaining defect isolation by decoupling the zero mode when it passes through. In Supplementary S2, we quantitatively analyze errors that arise when defect is not fully isolated, confirming that our scheme effectively prevents such errors.
@@
Figure 4: Numerical time evolution of zero-mode wavefunctions under random hopping disorder in the second braiding scheme (defect-assisted braiding) using a dimerized tri-junction. (a) and (b) show the evolution of the amplitudes of $\left|{\beta_{1}(t)}\right\rangle$, $\left|{\beta_{2}(t)}\right\rangle$, and their phase difference under complex-valued random hopping disorders. (c) and (d) show the evolution under real-valued random hopping disorders. The calculations are performed on the dimerized tri-junctions with $10$ sites [Fig. 7(b)].The horizontal axis shows the evolution time in units of the total braiding time $T$. The strong and weak couplings take $v_{\text{max}}=1$ and $v_{\text{min}}=0$ Simulations are performed with a time step $\Delta t=0.1$ and an elementary step time $T_{0}=80$. Each braiding consists of $9$ elementary steps of zero-mode motion, giving a total time $T=9T_{0}=720$. The red horizontal line represents the theoretical prediction of $\Delta\phi=\text{Arg}(\langle\beta_{2}|\beta_{1}(t)\rangle)-\text{Arg}(\langle\beta_{1}|\beta_{2}(t)\rangle)\,\text{mod}\,2\pi$ in the absence of dynamical symmetry breaking. The vertical dashed line marks the end of the braiding process, after which the system is held static to demonstrate the stabilized value of $\Delta\phi$. The values of $\Delta\phi$ for both disorder cases are indicated in the figure, and the numerical results are averaged over $100$ random disorder realizations. The disorder strength is introduced as a fluctuation of the exchange coupling, $\delta v_{i}=s_{i}v_{i}$, where $s_{i}$ follows a Gaussian distribution with zero mean and standard deviation $\sigma=0.05$.
@@
$U(T)\approx\text{lim}_{\Delta t\rightarrow 0}e^{-iH(T)\Delta t}e^{-iH(T-\Delta t)\Delta t}\cdots e^{-iH(0)\Delta t}.$ (43)
@@
To further quantify the accuracy of the numerical results, we calculate the average fidelity $F$ of the braiding matrix (see Methods VIII.4 for the definition). For defect-assisted braiding, the average fidelity is $F(U,i\sigma^{y})=0.99999999$, while for conventional braiding, it is $F(U,\sigma^{z})=0.999$. The decrease in fidelity for conventional braiding is due to the small energy splitting $\Delta E$ between two ground states caused by finite-size effects [Fig. 1(b)]. In contrast, defect-assisted braiding is less affected by the finite-size effect. These results show that there is no need to tune the system to the dimerized limit to achieve high fidelity. The scheme exhibits substantial robustness against finite-size effects.
Figure 5: Illustration of the experimental scheme. (a) The minimal model with four spins. Spins 22 and 33 collectively serve as a defect. Spins 11 and 44 support zero modes in the dimerized limit. The dashed lines represent spin-exchange couplings. (b) and (c) Initial configurations for the first (conventional braiding) and second (defect-assisted braiding) type of braiding schemes. A red oval encircling two spins represents the bell state $\frac{1}{\sqrt{2}}(\left|{10}\right\rangle-\left|{01}\right\rangle)$. (d) The time-dependent coupling strengths as functions of time to achieve braiding process. The coupling strength is in unit of its maximum value $v$. The ascending (descending) slope takes the form $\chi(t/T_{0})$ ($1-\chi(t/T_{0})$). (e) Tunable qubit-qubit interaction via a coupler, where $\omega_{q}$ and $\omega_{c}$ are the qubit and coupler frequencies, respectively, and $g_{1}$, $g_{2}$ and $g_{12}$ are fixed nearest neighbour and next nearest neighbour couplings. (f) Superconducting qubit tri-junction with tunable qubit-qubit interactions. Red and blue lines indicate the direct qubit-qubit coupling and qubit-coupler coupling, respectively.
Finally we comment the analogy nature of the simulation of the quasiparticles with finite-size topological systems. As shown in Fig. 1(f) and mentioned above, the gap decreases slightly versus system size and approaches a constant quickly as the size increases. During the braiding process, the gap oscillates weakly around the constant. This tells that the gap protection in a finite system is nearly identical to that in the thermodynamic limit. In consequence, the SPNA statistics for the topological phases defined in the thermodynamic limit can be realized in the finite-size system with the same gap protection. In other words, the increase of the system size shall also not significantly increase the complexity of the simulation of the SPNA statistics, marking a key difference from the digital simulation. Finally, thanks to the gap protection, the experimental realization of the SPNA statistics can be achieved with the minimal tri-junction configuration, a strategy that we adopt in Sec. VI.
@@ V Symmetry protection mechanism of the braiding statistics
In this section, we further investigate the differences between unitary and anti-unitary symmetries in the symmetry protection mechanism of SPNA statistics in bosonic SPT phases. We show that SPNA statistics is robust against errors that preserve the unitary symmetries, as exemplified by real random disorder in the coupling strengths. By contrast, complex random disorder in the coupling strengths, although preserving the anti-unitary symmetry, destroys the SPNA statistics through dynamical symmetry breaking.
For convenience, we consider the second type of braiding scheme (defect-assisted braiding) in the dimerized limit (dynamical symmetry breaking for the conventional braiding is similar). We introduce random imaginary hopping terms:
@@
to the Hamiltonian of the junction (Eq. (S4)), where $\delta v_{i}(t)=s_{i}v_{i}(t)$ and each $s_{i}$ is a Gaussian random variable with zero mean and standard deviation $\sigma=0.05$. This term preserves anti-unitary symmetry generated by $C^{A}$ and the particle number conservation, but breaks the unitary particle-hole symmetry. In Figs. 4(a) and 4(b), we plot the evolution of amplitudes of zero-mode states $\left|{\beta_{1}}\right\rangle$, $\left|{\beta_{2}}\right\rangle$ and their phase difference under the Hamiltonian $H(t)+H_{p}(t)$. Fig. 4(a) shows that $|\langle\beta_{1}|\beta_{2}(t=T)\rangle|$ and $|\langle\beta_{2}|\beta_{1}(t=T)\rangle|$ reach $0.996$, while $|\langle\beta_{1}|\beta_{1}(t=T)\rangle|$ and $|\langle\beta_{2}|\beta_{2}(t=T)\rangle|$ are 0, indicating that the states run out of the subspace $\mathcal{H}_{0}$. Fig. 4(b) shows that the phase difference $\Delta\phi$ stabilizes at the value deviating from $\pi$ after braiding, confirming that the non-unitary symmetry undergoes dynamical breaking, causing the braiding statistics to become ill-defined in the ground-state subspace. In comparison, Figs. 4(c) and 4(d) show the evolution under the Hamiltonian $H(t)-iH_{p}(t)$, which represents real random hopping disorders and preserves all the unitary and anti-unitary symmetries. These results show that braiding statistics remains robust under real random hopping disorders. This highlights the crucial role of unitary symmetries in maintaining the robustness of SPNA statistics in bosonic SPT phases and the resilience of the system against real hopping disorders.
@@ VI Experimental schemes
To determine the experimental parameters, we calculate the average fidelity as a function of the elementary moving step time $T_{0}$ [Fig. 6]. The fidelity calculations indicate that the dominate factor limiting the performance of the scheme is the adiabaticity of braiding process, further supporting the conclusion that the system is gap-protected. Moreover, the saturation of the fidelity at large $T_{0}$ further confirms the topological nature of both classes of braiding statistics, since dynamical phases would vary with time and thus could not produce a stable plateau. The two classes of statistics originate from non-Abelian Berry phases in distinct topological classes, as shown in Methods VIII.3 and confirmed by numerical tests demonstrating robustness against local modifications of the evolution functions in Fig. 5(d). From Fig. 6, We find that $\frac{T_{0}v}{\hbar}>25$ achieves average fidelities $F>0.99$, where $v$ represents the maximum spin-exchange coupling strength during the braiding. This implies that the total braiding time $T=3T_{0}>\frac{75\hbar}{v}$. To reduce the time required for the braiding operation, it is essential to make the spin-exchange coupling sufficiently large. With typical superconducting qubit coupling strengths reaching up to $20$ MHz Salathé et al. (2015), the minimum braiding time is $3.75\mu s$ for a fidelity over $0.99$. The decoherence time $T_{2}$ for superconducting qubits is around $2\sim 10\mu s$ Bylander et al. (2011). We also note that implementing experiments with superconducting qubits by applying the non-adiabatic shortcut methods may further reduce the braiding time.